# Audio-Language Models for Audio-Centric Tasks: A survey

### v1

@@ Abstract
Audio-Language Models (ALMs), which are trained on audio-text data, focus on the processing, understanding, and reasoning of sounds. Unlike traditional supervised learning approaches learning from predefined labels, ALMs utilize natural language as a supervision signal, which is more suitable for describing complex real-world audio recordings. ALMs demonstrate strong zero-shot capabilities and can be flexibly adapted to diverse downstream tasks. These strengths not only enhance the accuracy and generalization of audio processing tasks but also promote the development of models that more closely resemble human auditory perception and comprehension. Recent advances in ALMs have positioned them at the forefront of computer audition research, inspiring a surge of efforts to advance ALM technologies. Despite rapid progress in the field of ALMs, there is still a notable lack of systematic surveys that comprehensively organize and analyze developments. This deficiency not only limits researchers’ comprehensive understanding and evaluation of existing technologies but also hinders the rapid adoption and improvement of new methods. In this paper, we present a comprehensive review of ALMs with a focus on general audio tasks, aiming to fill this gap by providing a structured and holistic overview of ALMs. Specifically, we cover: (1) the background of computer audition and audio-language models; (2) the foundational aspects of ALMs, including prevalent network architectures, training objectives, and evaluation methods; (3) foundational pre-training and audio-language pre-training approaches; (4) task-specific fine-tuning, multi-task tuning and agent systems for downstream applications; (5) datasets and benchmarks; and (6) current challenges and future directions. Our review provides a clear technical roadmap for researchers to understand the development and future trends of existing technologies, offering valuable references for implementation in real-world scenarios.
@@ I Introduction
Enabling machines to hear like humans and process audio-centric tasks has long been a significant challenge [1]. Audio-Language Models (ALMs), which are trained on audio-text data, focus on the processing, understanding, and reasoning of sounds. This area is emerging as a prominent research field at the intersection of audio processing and Natural Language Processing. ALMs are not only applicable to basic audio tasks, such as audio classification [2], but also show great potential for more complicated scenarios. These include tasks such as audio-text retrieval [3], audio generation [4], automatic audio captioning [5], audio source separation [6], automatic speech translation [7], and audio chatbots [8].
@@
In contrast to audio representation learning based on labeled data for specific tasks, ALM can learn from more descriptive textual information, expanding the scope of supervision to include human-annotated captions and readily available titles and descriptions from web sources [9]. Natural language is well-suited for characterizing real-world audio, which frequently involves multiple overlapping sound events, thereby enabling models to learn their intrinsic relationships [10]. Furthermore, using natural language as supervision avoids the model’s reliance on task-specific predefined labels, enhancing the potential for models to generalize effectively to open-world scenarios.
@@
As large language models (LLMs) exhibit remarkable comprehension capabilities, researchers have explored their integration as guiding components within ALMs. However, pre-trained LLMs still face challenges in generalizing across a broad spectrum of downstream tasks [11], necessitating additional transfer steps such as post-training and collaboration with other foundational models. Within this research landscape, language provides a unified mechanism for constructing instances, enabling LLMs to undergo instruction tuning and in-context learning across diverse tasks. This approach bridges the gap between auditory information and language understanding, facilitating the alignment of multiple components within ALMs. Furthermore, language serves as a versatile human-machine interface, empowering users to instruct LLM agents to collaborate effectively with audio-language systems.
Despite the strong interest shown by the audio community in ALMs, there is still a lack of comprehensive surveys to review the current research status. Existing relevant reviews include speech-language models [12, 13], codec-based models [14], ALMs for specific tasks such as audio-text retrieval [15], automated audio captioning [16], speech-to-text translation [17], and audio-language datasets [18]. Here, we present the first comprehensive survey on ALMs, aiming to achieve an exhaustive coverage of the entire ALM research landscape from the perspective of model training. Additionally, we adopt a perspective centered on general audio-centric tasks that encompasses a diverse range of audio types to provide a more detailed reflection of the current state and development of computer audition. This survey method reflects mutual promotion and constraints among different research aspects from model to data, aids in systematically summarizes challenges and future directions, and serves as a guide for researchers and practitioners interested in ALM techniques, thereby facilitating further academic research and industrial applications in the field.
We first look at recent advances in ALM research and draw the timeline as shown in Fig.1. CLAP[2] is considered a significant milestone. Previous work includes some audio-caption datasets [19, 20, 21], which were initially used for automatic audio caption model training and also served as data foundations for ALMs, inspiring subsequent work. Since the introduction of pre-training and large-scale datasets [22], the advantages of ALMs have gradually gained attention. Recently, numerous new works have emerged, primarily reflecting the intertwined development between pre-training and downstream models. With increasing model research, recent studies have focused on the lack of unified evaluation standards and proposed various benchmarks. It shows a high correlation between datasets, pre-training, downstream models, and benchmark research in ALMs. Additionally, we observe that, driven by commercial applications, research interests have shifted more towards the speech domain. However, audio typically encompasses a variety of environmental events, including human voices, natural sounds, music rhythms, etc., which presents significant challenges to general audio modeling [23].
@@
Figure 1: A timeline of recent advances in audio-language models. Is is established mainly according to the release date (e.g., the submission date to arXiv) and some still working in progress. It highlights that datasets serve as the foundation for inspiring research in pre-training and downstream models. With the advancement of model research, recent studies have developed several benchmarks to promote comprehensive development in the field.
@@
In the subsequent sections of this paper, we first introduce the background of audio-language pre-training and transfer paradigm (Section II). We then describe the foundations of ALMs, including model architecture, training objectives, and evaluation methods (Section III). Following this, we review the topics of representation pre-training (Section IV), downstream transfer (Section V), and related data (Section VI). Building on these foundations, we discuss the challenges and future research directions (Section VII), before concluding the paper (Section VIII).
@@ II Background
This section begins by discussing the development of computer audition paradigms, with a particular focus on how ALMs are trained and transfer for downstream, as well as the reasons for the shift towards the audio-language paradigm. We then introduce the training stages and establish a research landscape for ALMs, providing a structured basis for the comprehensive review in the following sections.
@@ II-A Pre-training and Transfer Paradigm
The pre-training and transfer paradigm involves initially training on large-scale public datasets to get robust representations, and then applying knowledge gained from one context to another to enhance the performance on downstream tasks. This approach accelerates supervised learning on downstream tasks.
However, as this paradigm evolves, two challenges emerge. First, models may overfit by exploiting simple label mappings, achieving high performance on specific tasks without truly understanding the underlying audio content [24], leading to poor generalization to new data. Second, the high cost of manual annotation exacerbates the difficulty of obtaining limited labeled datasets for learning audio representation [25].
To address these challenges, ALMs have been proposed to learn audio concepts through natural language supervision [2]. Firstly, this form of supervision provides more details about the audio, enabling models to understand the meanings and make decisions accordingly like a human. For example, natural language can describe the temporal order of multiple events using words such as ‘simultaneous,’ ‘before,’ and ‘after’ [26], better reflecting the complex composition of audio compared to predefined labels and helping models learn their intrinsic relationships [10]. Additionally, audio-text data is easier to obtain than well-defined labeled datasets, effectively expanding the scale of datasets. For instance, we can use ‘dog’ or ‘barking’ to label a dog barking, but inconsistencies among multiple annotators make it difficult to create a perfectly accurate audio dataset. While ALMs are able to leverage the natural language processing capabilities of pre-trained models to extract similar semantic features from different forms of descriptions. Besides human-annotated captions and translations, titles and descriptions related to audio found abundantly on the web can also serve as sources of text annotation.
@@ II-B Audio-Language Training Stages
As data and model sizes grow, the training strategies for ALMs become more intricate. From the viewpoints of representation learning and downstream task application, we first categorize the training stages aimed at enhancing task-independent audio representations as falling within the scope of pre-training, while fine-tuning and cooperating before the model is applied to downstream tasks are defined as part of the transfer process.
ALMs pre-training can be further divided into multiple stages, typically including the pre-training of foundational models, followed by audio-language pre-training on paired data. Some may also involve further training on a broader range of data and tasks.
Although ALMs have achieved strong zero-shot capabilities in audio retrieval, transfer remains an important stage for applying models to downstream tasks. Task-specific fine-tuning is one of the most widely used methods. It involves supervised fine-tuning of pre-trained models on downstream task datasets and may require the addition of some adaptive modules. Another category of methods includes transferring simultaneously on multiple tasks to make the model more universal or gain from multi-task knowledge sharing. Unlike task-specific fine-tuning, which focuses directly on task performance, instruction tuning and in-context learning aim to enhance (or unlock) the LLM’s ability to follow human instructions. Essentially, it fine-tunes ALMs with a set of formatted instances in natural language form [27], thus helping the model generalize to various downstream tasks. Multi-task transfer can also be achieved by cooperating multiple models to form an agent system.
@@ II-C Research Landscape
Based on current research and our definition of audio-language training stages, we construct a research landscape for ALMs, as shown in Fig. 2. From the training dimension, ALMs are divided into pre-training and transfer. ALMs achieve multimodal perception by integrating pre-trained audio and language models, then undergo further pre-training on extensive audio-text data. Transfer is crucial for combining these models with other networks and applying them to various downstream tasks. Data is an essential element for model training and evaluation. Different types of datasets can be utilized at various stages of training, and benchmarks provide unified and comprehensive standards for model evaluation, playing an important role in optimizing the models. Therefore, research on ALMs can be developed in three fields: (a) pre-training for representation learning, (b) downstream transfer, and (c) datasets and benchmarks.
Figure 2: Research landscape for audio-language models. From the perspective of model training, (a) audio-language representation requires pre-training (Sec. IV), (b) transfer to downstream application through task-specific fine-tuning or instruction tuning (Sec. V), (c) data is the foundation for model training, and they can be divided into labeled audio datasets, audio-text paired datasets, and audio question answering datasets (Sec. VI).
Within the scope of the research landscape, we designed a review outline as shown in Fig. 3. We first provide an overview of the foundation on ALMs, thereby comprehensively reviewing related work from three research fields. According the progress across areas, we systematically propose the challenges and future directions for ALMs.
@@ III Foundations
In this section, we will introduce the general foundations of ALMs, including commonly-used architectures, training objectives, and evaluation methods.
@@ III-A ALM Architectures
Audio-language models and systems typically comprising audio and language encoders, and may include other multimodal alignment mechanisms and language models. As shown in Fig. 4, current ALMs can generally be divided into four types: Two Towers, Two Heads, One Head and Cooperated Systems.
@@ III-A1 Two Towers
The basic form of ALMs, with one encoder and a projector for each modality, embeddings will be aligned in a joint space. Among them, the most prominent landmark pretraining research is Contrastive Language-Audio Pretraining (CLAP), which incorporates a contrastive learning framework to bring audio and text descriptions into a joint multimodal space, learning the mapping relationship between the two modalities [2]. Furthermore, based on the concept of modality alignment, mechanisms can be added between two independent encoders to facilitate communication, with the aim of achieving early-stage modality fusion during the representation phase [28].
@@ III-A2 Two Heads
A mainstream form that utilizes one encoder and a projector for each modality, with a language model on top. Here, ‘Head’ refers to a network that unifies a certain modal representation space into a unified space [29, 30, 31]. Language modeling has first been proven to possess strong capabilities in semantic feature extraction within the field of speech [32], making it a natural design choice to incorporate language models into ALMs. With the development of large language models, many works have utilized LLMs as the backbone for ALM inference, expanding the perceptual modalities of large language models and leveraging their emergent understanding capabilities. This has led to classic works such as SpeechGPT [8], Pengi [1], and Qwen-Audio [33], making Two Heads a unified architecture of Large Audio-Language Models. In this structure, modality fusion can also be promoted through communication mechanisms between encoders [34]. It is important to note that in some works, text inputs may only undergo tokenization without the need for a dedicated text encoder, and these models can be considered under a special type of Two Heads framework.
@@ III-A3 One Head
A unified multimodal input form that uses one encoder to handle two different modalities simultaneously, with a language model on top. In the vision community, a line of work has conducted research on the One Head architecture based on the view that the same multimodal processing module can achieve better alignment. That is, using a unified space to represent two modalities. However, there are relatively few related studies in audio-language [35].
@@ III-A4 Cooperated Systems
This system employs an LLM as a planning agent and comprises various model types mentioned above. Its design facilitates the selection and utilization of each model’s inherent complementary strengths, tailored to downstream task requirements. Through the collaboration of these diverse models, the system can tackle a wider array of complex tasks compared to a solitary model alone [36].
Figure 4: Typical architectures of audio-language models. (a) Two Towers, with one encoder and a projector for each modality, embeddings will be aligned in a joint space. (b) Two Heads, adds language model on top. (c) One Head, with one unified encoder and a language model. (d) Cooperated Systems, utilize LLMs as agents to cooperate several models.
@@ III-B Training Objectives
Training objectives are used to guide model learning during pre-training and transfer. As shown in Fig. 5(a), pre-training contrastive, generative, or discriminative objectives guide the model to learn pretext tasks on audio, text, or audio-text paired data, aiming to learn audio semantic features and audio-language correlations. As illustrated in Fig. 5(b), task-specific fine-tuning as a commonly adopted transfer method, employs either generative or discriminative objectives depending on the context. Another line of transfer methods with generative language models in Fig. 5(c) aims to improve unlock pre-training models’ generalization ability on downstream tasks through standard language modeling objectives. Note that the above training objectives can be used in combination.
@@ III-B1 Contrastive Objectives
It is the most commonly used type of training objective in audio-language pre-training, which aims to train the model to bring positive sample pairs closer together and push negative sample pairs further apart within a shared embedding space for audio and text, thereby learning the audio-language correlations and obtaining distinguishable representations between audio samples. The most widely implemented approach for this category of objective is using a symmetric audio-text infoNCE [37] loss function to measure the similarity between audio and text embeddings. Let the $i-th$ sample pair be ${x_{i},t_{i}}$. Given an audio encoder $h_{a}(\cdot)$ and a text encoder $h_{t}(\cdot)$, the embedding vectors for the audio sample $x_{i}$ and its corresponding caption $t_{i}$ can be represented as:
$z_{i}^{a}=h_{a}(x_{i})$ (1)
$z_{i}^{t}=h_{t}(t_{i})$ (2)
The similarity between audio and text embeddings is calculated using the dot product. The infoNCE loss for the audio dimension, $l_{a}$, is defined as the average of a normalized function measuring the similarity of different texts to the same audio query. Similarly, the contrastive loss for the text dimension, $l_{t}$, measures the similarity of different audios to the same text query. For a batch with $B$ audio-text pairs, we have:
@@
where $\tau$ represents a temperature parameter used to scale the range of logits. When setting the contrastive objective to be completely symmetrical, the total loss for the audio-text pairs in one batch can be defined as:

### v2

# Audio-Language Models for Audio-Centric Tasks: A Systematic Survey

@@ Abstract
Audio-Language Models (ALMs), trained on paired audio-text data, are designed to process, understand, and reason about audio-centric multimodal content. Unlike traditional supervised approaches that use predefined labels, ALMs leverage natural language supervision to better handle complex real-world audio scenes with multiple overlapping events. While demonstrating impressive zero-shot and task generalization capabilities, there is still a notable lack of systematic surveys that comprehensively organize and analyze developments. In this paper, we present the first systematic review of ALMs with three main contributions: (1) comprehensive coverage of ALM works across speech, music, and sound from a general audio perspective; (2) a unified taxonomy of ALM foundations, including model architectures and training objectives; (3) establishment of a research landscape capturing mutual promotion and constraints among different research aspects, aiding in summarizing evaluations, limitations, concerns and promising directions. Our review contributes to helping researchers understand the development of existing technologies and future trends, while also providing valuable references for implementation in practical applications.
@@ I Introduction
Enabling machines to hear like humans and process audio-centric tasks has long been a significant challenge [deshmukh2023pengi]. Audio-Language Models (ALMs), which are trained on audio-text data, focus on the processing, understanding, and reasoning of audio. This area is emerging as a prominent research field at the intersection of audio processing and natural language processing. ALMs are not only applicable to basic audio tasks, such as audio classification [elizalde2023clap] and sound event detection, but also show great potential for more complicated scenarios. These include tasks such as audio generation [liu2023audioldm], automatic audio captioning [kim2024enclap], audio source separation [liu2022_lass], and audio chatbots [zhang2023speechgpt].
@@
In contrast to audio representation learning based on labeled data for specific tasks, ALM learns from more descriptive textual information, expanding supervision to include human-annotated captions and web sources like titles and descriptions [laion_ai_audio_dataset]. Natural language effectively characterizes real-world audio, which often involves multiple overlapping sound events, enabling models to learn their intrinsic relationships [wu2019audio]. Moreover, using natural language as supervision eliminates the model’s dependency on predefined task-specific labels, enhancing its potential to generalize in open-world scenarios.
@@
Leveraging the strong comprehension capabilities of large language models (LLMs), researchers have integrated them as reasoning components into ALMs, sparking the emergence of Large Audio-Language Models (LALMs). However, pretrained LLMs still face generalization challenges across diverse downstream tasks [zhao2023LLMsurvey], requiring additional transfer steps such as post-training and collaboration with other foundation models. Within this research landscape, language offers a unified mechanism for constructing instances, enabling LLMs to undergo instruction tuning and in-context learning across various tasks. This bridges auditory information with language understanding while aligning multiple ALM components. Furthermore, language serves as a versatile human-machine interface, allowing users to direct LLM agents in coordinating effectively with audio-language systems.
Despite considerable interest in audio-language models (ALMs), no comprehensive survey currently captures the full scope of this rapidly evolving field. Existing relevant reviews include speech-language models [cui2024speechlanguage, ji2024wavchat, arora2025landscape, Peng2025SLLMsurvey], codec-based models [wang2023codec], ALMs for specific tasks such as audio-text retrieval [koepke2022benchmarks], automated audio captioning [xu2023aac_survey], speech-to-text translation [xu2023S2TT], audio-language datasets [wijngaard2024ald_survey], and evaluations of LALMs [yang2025evaluationsurvey]. While these reviews offer valuable insights into ALM subdomains, their narrow focus fails to capture broader developments and cross-domain synergies. This fragmentation motivates our work: the first comprehensive ALM survey. Our three key contributions include: (1) adopting a general audio-centric perspective to cover diverse works on human voices, natural sounds, and music, offering a complete picture of computer audition progress; (2) providing a unified taxonomy of ALM foundations through systematic review, including architectures and training objectives; (3) establishing a training-oriented research landscape that captures mutual promotion and constraints among different research aspects (from model to data), aiding in systematic summarization of evluations, limitations, concerns and future directions.
To contextualize our contributions, Fig. 1 illustrates the evolution of ALMs, highlighting key developments across research aspects. Early audio-caption datasets [kim2019audiocaps, drossos2020clotho, lipping2022clotho] enabled significant pre-training advances like CLAP [elizalde2023clap]. Subsequent large-scale datasets [wu2023large] demonstrated ALMs’ generalization capacity, accelerating progress in downstream tasks and benchmarks. This progression aligns with our research landscape (Fig. 2), where pre-training, transfer learning, datasets, and benchmarks interact synergistically. We also note growing emphasis on speech applications, though general audio modeling remains challenging due to diverse sound events [chen2023beats].
@@
Figure 1: A timeline of recent advances in audio-language models, organized primarily by release dates (e.g., arXiv submission) with some works still in progress. It highlights that datasets serve as the foundation for pre-training and downstream research. As model research evolves, recent studies have developed multiple benchmarks to drive comprehensive field development.
@@
In the following sections, we first present the background of audio-language pre-training and transfer learning (Section II), followed by foundations of ALMs (III). We then examine representation pre-training (IV), downstream transfer (V), along with related datasets (VI) and evaluations (VII). Based on this review, we discuss key limitations, concerns, and future directions (VIII, IX), and conclude the paper (X).
@@ II-A Pre-training and Transfer Paradigm
The pre-training and transfer paradigm involves training on large-scale public datasets to learn robust representations, which are then adapted to downstream tasks. While this approach accelerates supervised learning, it faces two key challenges. First, models may overfit by relying on superficial label correlations rather than learning meaningful audio content, limiting generalization [gong2024ltu]. Second, the high cost of manual annotation constrains the scale of labeled data, hindering representation learning [sun2024AAA]. To address these, ALMs leverage natural language as supervision. This offers richer audio descriptions, such as temporal relationships between events (e.g., “simultaneous” or “before”), enabling models to better understand complex auditory scenes [ghosh2024compa, wu2019audio]. Additionally, audio-text data is more scalable. While a dog barking might be variably described as ”dog” or ”barking,” leading to annotation inconsistencies, ALMs are able to extract consistent semantic features from diverse descriptions. Beyond human-annotated texts, web-sourced titles and descriptions further expand available annotations.
@@ II-B Audio-Language Training Stages
As data volumes and model complexity grow, ALM training strategies have become increasingly sophisticated. We distinguish two main stages from representation learning and downstream application: pre-training develops task-agnostic audio representations, while transfer learning adapts models through fine-tuning and cooperative methods. Pre-training typically involves multiple phases: initial foundation model training, audio-language alignment on paired data, and potentially extended training with diverse datasets and tasks. Despite strong zero-shot retrieval capabilities, transfer learning remains essential for most downstream applications. Task-specific fine-tuning represents the most common approach, often incorporating adaptive modules. Multi-task transfer leverages shared knowledge across tasks through joint training. Beyond task-specific optimization, instruction tuning and in-context learning enhance LLMs’ instruction-following abilities through natural language prompts [wei2022instruction], thereby improving generalization. Multi-task transfer can also be achieved via collaborative agent systems.
@@ II-C Research Landscape
To systematically survey ALMs, we introduce the research landscape in Fig. 2, which captures interconnections across research areas through two core stages: (a) Pre-training, which integrates pre-trained audio and language models and aligns them using large-scale audio-text data to achieve multimodal perception capabilities; and (b) Transfer, which adapts and combines pre-trained models with other components to support diverse downstream applications. The development of both stages relies on (c) Data and Benchmarks, where datasets provide resources for model training, while benchmarks establish unified standards for evaluation and optimization.
Guided by this landscape, our review outline in Fig. 3 follows a systematic progression: analyzing ALM foundations, reviewing advances across core stages, examining how data and benchmarks drive progress, and culminating in a synthesis of cross-domain challenges and future directions.
Figure 2: Research Landscape: (a) model pre-training for audio-language representations (Sec. IV), (b) model transfer via task-specific fine-tuning or multi-task cooperation (Sec. V), (c) foundational data including audio-text pairs, QA datasets, and evaluation benchmarks (Sec. VI).
@@ III-A ALM Architectures
ALMs typically comprise audio and language encoders, and may include multimodal alignment mechanisms and language modeling networks. As shown in Fig. 4, we categorize ALMs into four main types: Two Towers, Two Heads, One Head, and Cooperated Systems. These architectures, defined by distinct data flow patterns embody different multimodal philosophies, and directly influence deployment and performance.
@@ III-A1 Two Towers
This basic architecture employs separate encoders and projectors for each modality, aligning embeddings in a joint space. Exemplified by the landmark CLAP model, it uses contrastive learning to map audio and text into a shared multimodal representation [elizalde2023clap]. This late-interaction design yields low inference complexity, with generalization strengths in cross-modal retrieval and zero-shot classification. By shifting online inference to similarity computation through pre-computed cached representations, Two-Tower architectures achieve low-latency deployment for large-scale retrieval. While typically requiring hundreds of millions of parameters, memory can be reduced via knowledge distillation [paissan2024tinyclap]. Alignment mechanisms can be integrated between encoders to enhance representation learning [li2021ctal].
@@ III-A2 Two Heads
This mainstream architecture adds a language model atop separate modality encoders and projectors. The ’Head’ projects modal representations into a unified space [jang2023oner]. Language modeling enhances speech tasks [wang2023VALLE], while evolving LLMs serve as reasoning backbones [deshmukh2023pengi, chu2023qwenaudio]. Modality fusion can be enhanced through encoder alignment mechanisms [zhao2024mint], and text may be directly tokenized without a dedicated encoder [das2024speechverse]. This intermediate-fusion architecture, inference latency and memory requirements are dominated by the LM, varying with model scale and generation mode (e.g. autoregressive/non-autoregressive). Internal variation across different LM choices is substantial. Initialization with powerful LMs aids convergence and generalization via knowledge transfer [kaplan2020scaling], though stability depends on module selection and fine-tuning strategies.
@@ III-A3 One Head
This unified architecture processes both modalities through a single encoder before decoding. While vision research shows shared multimodal processing improves alignment [jang2023oner], audio-language applications remain limited [sachidananda2022calm]. This early-fusion approach has primarily explored using joint representations for downstream task fine-tuning [sachidananda2022calm], with no existing work incorporating LLMs for decoding. Inference latency in such models primarily comes from the fusion module’s forward computation. Theoretically, a unimodal encoder has the potential to achieve faster inference efficiency than Two Heads architectures, but faces steeper convergence challenges from joint cross-modal optimization in a unified parameter space [xia2023unicode].
@@ III-A4 Cooperated Systems
This system employs an LLM as a planning agent and comprises various model types mentioned above. It selects and facilitates different models based on their complementary strengths. Through the collaboration of these diverse models, the system can tackle a wider array of complex tasks compared to a solitary model alone [li2024survey]. Their memory usage varies with the loaded foundation models, and inference latency depends on execution paths. They face unique convergence challenges from the interplay between component performance and policy learning, and achieve distinctive generalization through modular capability composition.
Figure 4: Typical architectures of audio-language models. (a) Two Towers, with one encoder and projector for each modality, embeddings will be aligned in a joint space. (b) Two Heads, adds language model on top. (c) One Head, with a unimoal encoder and decoder. (d) Cooperated Systems, utilize LLMs as agents to cooperate several models.
@@ III-B Training Objectives
Training objectives guide model learning during pre-training and transfer. As shown in Fig. 5(a), pre-training objectives include contrastive, generative, and discriminative losses that learn pretext tasks on audio, text, or audio-text pairs to capture semantic features and cross-modal correlations. Fig. 5(b) illustrates task-specific fine-tuning using generative or discriminative objectives, while Fig. 5(c) shows transfer via generative language modeling to enhance generalization. Note that the choice of training objective depends on the specific requirements of the learning stage and application scenario, and the above training objectives can be used in combination. For detailed ablation studies on objective combinations, see works like MINT [zhao2024mint, yeh2023flap].
@@ III-B1 Contrastive Objectives
Contrastive learning, widely used in audio-language pre-training, brings positive audio-text pairs closer while pushing negatives apart in a shared embedding space, thereby learning the audio-language correlations and obtaining distinguishable representations between audio samples [elizalde2023clap]. The most widely implemented approach for this category of objective is using a symmetric audio-text infoNCE [oord2018CPC] loss function to measure the similarity between audio and text embeddings. For a batch of $B$ pairs $\{x_{i},t_{i}\}$, with audio encoder $h_{a}(\cdot)$ and text encoder $h_{t}(\cdot)$, embeddings are:
$z_{i}^{a}=h_{a}(x_{i}),\quad z_{i}^{t}=h_{t}(t_{i})$ (1)
The similarity between audio and text embeddings is computed via dot product. For a batch of $B$ audio-text pairs, the loss for the audio dimension $l_{i}^{a}$ measures how well the audio query $z_{i}^{a}$ matches its corresponding text $z_{i}^{t}$ against all other texts in the batch. Similarly, the text dimension loss $l_{i}^{t}$ assesses the text query $z_{i}^{t}$ against all other audios:
@@
Here, $\tau$ is a temperature parameter scaling the logits. The symmetric contrastive loss combines both directions:
@@
Theoretically, contrastive loss is non-decomposable, since the loss for a positive pair couples with all negatives in the batch. This implies that a finite batch size limits the negative pool to a sparse, biased subset of the true distribution, introducing systematic gradient bias and suboptimal convergence [chen2022gradientbias]. While ALMs commonly adopt large batches (e.g., 2304 in FLAP [yeh2023flap]) and have explored hard negative mining [li2024MGACLAP] to mitigate sampling bias, the problem itself remains underexamined in the community. Large-batch training improves performance but incurs high computational costs. More principled solutions from contrastive learning, such as momentum encoders [he2020moco] and debiased techniques based on importance weighting [chuang2020debiased], decouple negative pool size from batch constraints or correct sampling bias through weighting, enabling more accurate estimation of the true contrastive loss. We highlight these as promising yet unexplored directions for future contrastive ALM research.